LHC 2011 - latest news

Week 46: Machine coordinators: Ralph Assmann, Jan Uythoven

Goals of the week: Establish high intensity Pb-Pb physics. Test p-Pb physics.




Link to report from morning meeting

Monday 21st Nov

  • 08:45 Cryogenics S34 back. Problems access doors.
  • 10:30 Injecting pilot.

8:30 meeting

Sunday 20th Nov

  • 02:43 Stable beams Fill #2308.
  • 09:27 Beam dump by operator
  • 12:53 Stable beams Fill #2309
    Initial  lumi (Atlas) 4.25e26 cm-2s-1
  • 14:30 Life time of beams suddenly going down, recovered in few minutes
  • 17:29 Beam dump, preventive dump, loss cryo maintain S34.
    Cold compressor trip Point 4.
  • 19:00 Intervention for MKD switch replacement to be done now, teams preparing Also EPC imtervention on RCBCV10.R3B1 is set for 8pm

9:00 meeting

Saturday 19th Nov

  • 05:15 Beam dump by operator
  • 08:10 Stable beams Fill #2305.
  • 11:08 Lost QPS ok  in B11.L1 (location of ion losses). QPS piquet called.
  • 14:24 Beam dump. XPOC Beam1 blocked due to BLM: BLM on TCDSA is reproducibly just over threshold.
  • MB11.L1 QPS power cycle.
  • 16:33 Injection. Missing batch/low intensity batch when injecting. Dump and re-inject.
  • 19:56 Adjust.
  • 20:04 Beams dumped. low filament current on RF line 2B1 caused module M1B1 to trip.
  • 21:53 RF intervention finished after reset of regulator on klystron.
  • 23:45 RF filament current error on cavity 2B1 at flat top. Keep cavity 2B1 off. B1 RF voltage reduced from 14MV to 12.2MV.

9:00 meeting

Friday 18th Nov

  • 00:35 Stable beams. Fill #2302.
  • 06:07 Beam dump by RF temperature interlock (HOM 4B2). Total delivered Pb-Pb lumi in 2011 reached 19.6 ub-1 in ATLAS.
  • 09:34. Stable beams. Fill #2303.
  • 14:48 Beams dumped.
    Spurious opening of the quench loop on the RQF in S12
  • 16:35 QPS access for the switch on RDC.A12B2 . ALICE also in.
  • 18:30 end of access
  • 22:41 Stable beams Fill #2304

8:30 meeting

Thursday 17th Nov

  • 00:21 Stable beams
  • 09:00 Dumped beams, refill for Pb-Pb physics. No more protons in 2011…
  • Increase of flat top total voltage to 14 MV. Feed-forward on the H orbit CODs in the ramp and squeeze. New version of the FBCT.
  • Release of SIS. The only change concerns the injection bucket check. The logic now allows injection if there is a lone single bunch in the machine beyond the point of the incoming injection.
  • 11:00 Problem on dump kicker B1, one swithc is dying and may need to be exchanged. It has been masked for the time being to allow to go in physics.
  • 12:56 Stable beams. Fill #2301.
  • 13:47 Changed the settings on the acquisitions for the BQMLHC so that the pattern for ring 2 is not detected incorrectly at the flat top (so far at times the last batch was detected 1 bucket aside). The acquisition was always ok at the flat bottom, where it is vital...
  • 18:46 change the target length for the blow up from 1.05E-9 to 1.25e-9 for both beams. We will have longer bunches for the next ramp.
  • 19:22 Ramp down undulators for BSRT tests.
  • 19:59 Adjust mode: end of stable beams. 4.2-4.8ub-1 delivered per experiment. Test of lumi scan software changes for 25min. Some bug found -> not yet deployed.
  • 20:30 Beam dump. CMS access during ramp down and cycle.
  • 21:08 Undulator is back on.
  • 23:41 Ramp.

8:30 meeting

Wednesday 16th Nov

  • 00:05 Preventive beam dump due to cryo problem in 34. Fault on the power converter for the cold compressor. Collected 4-4.5 mb-1 per experiment in fill #2297.
  • 09:15 Cryo conditions back
  • 12:20 PS Septum problem - no beams available for a while...
  • Decide to go back to Pb-Pb for the night
  • 20:45 Injecting Pb pilots

8:30 meeting

LHCb gas injection

Tuesday 15th Nov

  • 00:20 Start ramp
  • 01:30 Stable beams, fill #2294 170 x 170 bunches. Lumi up to 1.5e26.
  • 08:25 Beams dumped due to RF vacuum interlock
  • 09:53 Triplet R2 power converter problem, access required
  • 11:50 Access required for the RF vacuum point 4.
  • 13:50 Start pre-cycle
  • 16:20 Injection
  • 17:15 Filled 358 on 358 bunches. Nominal ion intensity.
  • 17:23 Ramp.
  • 17:43 Squeeze.
  • 18:22 Stable beams. Fill #2297. Peak luminosity: 3.5e26 cm-2 s-1. 1.8MJ of stored energy per ion beam.
  • 22:29 G. Schneider is going to start the injection of gas in LHCb.

8:30 meeting 

Monday 14th Nov

  • 10:00 PS vacuum about recovered
  • 11:30 Injecting probes
  • Several problems on the injector chain in the afternoon and evening plus some time for setting up injection in the LHC

8:30 meeting